OpenStudy (anonymous):

If 0° < x° < 90° and sin x = 1/2 , then cos x = ?

myininaya (myininaya):



myininaya (myininaya):

\[\cos(x)=\frac{adj}{hyp}=\frac{adj}{2}\]

OpenStudy (anonymous):